### v4

# Shuchang Zhang, Hui Zhang, Hongxia Wang

Abstract
Recently, the degenerate preconditioned proximal point (PPP) method provides a unified and flexible framework for designing and analyzing operator-splitting algorithms such as Douglas-Rachford (DR). However, the degenerate PPP method exhibits weak convergence in the infinite-dimensional Hilbert space and lacks accelerated variants. To address these issues, we propose a Halpern-type PPP (HPPP) algorithm, which leverages the strong convergence and acceleration properties of Halpern’s iteration method. Moreover, we propose a novel algorithm for image restoration by combining HPPP with denoiser priors such as Plug-and-Play (PnP) prior, which can be viewed as an accelerated PnP method. Finally, numerical experiments including several toy examples and image restoration validate the effectiveness of our proposed algorithms.
keywords:
Halpern iteration, Preconditioned proximal point algorithms, Plug-and-Play Prior, Regularization by Denoising, Image restoration
1 Introduction
Image restoration (IR) problems, including image deblurring, super-resolution, and inpainting, can be formulated as the following optimization problem [16, 27]:
(1) $\min_{\mathbf{x}\in\mathcal{X}}\lambda f(\mathbf{x})+g(\mathbf{K}\mathbf{x}),$
where $f:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ and $g:\mathcal{Y}\to\mathbb{R}\cup\{+\infty\}$ are convex, lower semicontinuous functions, $\mathbf{K}:\mathcal{X}\to\mathcal{Y}$ is a bounded linear operator, and $\lambda>0$ is a balance parameter. Both $\mathcal{X}$ and $\mathcal{Y}$ are real Hilbert spaces. The first term $f$ represents the data fidelity, while the second term $g$ serves as a regularization such as TV [48] to mitigate the ill-posedness of IR problems.
By the first-order optimality condition, the convex optimization problem (1) is equivalent to the following inclusion problem:
[…]
where $\partial f(\mathbf{x})$ and $\partial g(\mathbf{x})$ are the subdifferentials of $f$ and $g$ at $\mathbf{x}$, respectively [5, Chapter 3]. Following [27, 9], by introducing an auxiliary variable $\mathbf{y}\in\partial g(\mathbf{K}\mathbf{x})$, we reformulate (2) as:
[…]
where $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&(\partial g)^{-1}\end{pmatrix}$, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, and $\mathcal{H}=\mathcal{X}\times\mathcal{Y}$. The problem (3) is common in modern optimization and variational analysis [2, 3]. When $\mathcal{A}$ is maximal monotone, the resolvent $J_{\mathcal{A}}=(I+\mathcal{A})^{-1}$ is nonexpansive with a full domain, as established by the Minty surjectivity theorem [39]. The proximal point iteration $\mathbf{u}^{k+1}=(I+\mathcal{A})^{-1}\mathbf{u}^{k}$ is used to solve (3) and converges weakly [46]. However, it is difficult to compute the operator $(I+\mathcal{A})^{-1}$, and hence splitting methods have been developed to address this issue. The well-known Douglas-Rachford splitting (DRS) [23] decomposes $\mathcal{A}$ into the sum of two maximal monotone operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, for which $J_{\mathcal{A}_{1}}$ and $J_{\mathcal{A}_{2}}$ are easier to obtain. Another way to solve (1) is to transform it into the saddle-point problem [16, 43], i.e.,
(4) $\min_{\mathbf{x}\in\mathcal{X}}\max_{\mathbf{y}\in\mathcal{Y}}\left\langle\mathbf{K}\mathbf{x},\mathbf{y}\right\rangle+\lambda f(\mathbf{x})-g^{*}(\mathbf{y}),$
where $g^{*}:\mathcal{K}\to\mathbb{R}\cup\{+\infty\}$ is the conjugate of $g$. In this direction, many primal-dual methods, including the well-known Chambolle-Pock (CP) splitting method [16, 43] and the Primal-Dual Hybrid Gradient (PDHG) method [27], are designed and studied extensively. In particular, He and Yuan [27] analyzed the PDHG method from a PPP standpoint with a positive definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$. Recently, Bredies et al. [11, 12, 9] developed a unified degenerate PPP algorithmic framework with a positive semi-definite $\mathcal{M}$. By choosing appropriate preconditioners, this framework could cover the DR and CP splitting algorithms [9].
It is crucial to study algorithms that exhibit both strong convergence and acceleration [30, 8, 34, 3, 52, 17, 38], such as strongly convergent proximal point methods [30] and forward-reflected-backward splitting algorithms [34]. For example, Bauschke et al. demonstrated strong convergence of degenerate PPP under the special case where $\mathcal{A}$ is linear [3]. In such cases, the weak limit of the PPP sequence corresponds to the $\mathcal{M}$-projection of the initial point. Moreover, Sun et al. [52] proposed an accelerated preconditioned Alternating Direction Method of Multipliers (pADMM) by leveraging the degenerate PPP method [9] and the fast Krasnosel’skiĭ-Mann (KM) iteration [6]. Similarly, Chen et al. introduced an accelerated HPR-LP solver, which implements a Halpern Peaceman-Rachford method enhanced with semi-proximal terms for efficiently solving linear programming (LP) problems [17], and the restarted Halpern PDHG (rHPDHG) achieves an accelerated linear convergence rate [38]. A shortcoming of forward-backward and DRS algorithms employing KM iteration is that their iterates converge only in the weak topology [8]. Consequently, the degenerate PPP method based on KM iteration typically exhibits weak convergence in Hilbert spaces. Meanwhile, it is unclear whether the degenerate PPP method can be accelerated via simple modification.
The classic Halpern iteration [26] offers the advantage of strong convergence over the KM iteration in infinite-dimensional Hilbert spaces, with the limit identified as the metric projection of the anchor onto the fixed point set [29]. Then the Halpern iteration is also known as an implicitly regularized method [22]. Due to this implicit regularity, the degenerate PPP method incorporating Halpern iteration can obtain a unique solution, yielding stable recovery results for IR problems. Achieving stable reconstruction for ill-posed inverse problems is important [24]. Moreover, beyond the implicit regularity, Halpern iteration can also accelerate convergence rate in terms of operator residual norm [55], which is widely utilized in machine learning [22, 41, 29]. Notably, the work [41] has demonstrated that PDHG with restarted Halpern iteration achieves a faster convergence rate for function values in CT image reconstruction. Based on the Halpern iteration, the Halpern-type preconditioned proximal point algorithm called HPPP (5) is thus proposed to overcome the two limitations of the degenerate PPP method,
(5) $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k},$
where $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$ are the anchor point and the initial point, respectively, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ is a sequence in $[0,1]$ such that $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty,\lim_{k\to\infty}\mu_{k}=0$.
PnP (Plug-and-Play) methods that combine splitting algorithms with denoiser priors have been widely applied in practical problems [56, 51, 1, 35, 59] and have achieved state-of-the-art performance in inverse imaging tasks [65, 33, 53]. Buzzard et al. provided a Consensus Equilibrium interpretation on denoiser priors [14]. Romano et al. introduced RED (Regularization by Denoising) [47], whose gradient exactly corresponds to the denoising residual, thereby yielding a clear objective function [20] that can be exploited in first-order optimization. Powerful denoisers (e.g., DnCNN [66]) typically do not meet the conditions of RED [45]. To address this limitation, Reehorst and Schniter introduced the score-matching by denoising (SMD) perspective to interpret RED [45]. Based on the fixed-point projection, Cohen et al. proposed the RED-PRO model [20] as a bridge between RED and PnP, while the hybrid steepest descent (HSD) method [63, 64] was employed to solve the resulting model. Meanwhile, Ryu et al. proved the convergence of PnP-FBS and PnP-ADMM using the Banach contraction principle under the assumption that the data term $f$ is strongly convex and that the residual is nonexpansive [49]. However, enforcing strong convexity on $f$ precludes many IR tasks [33]. Furthermore, Cohen et al. parameterized denoisers via the gradients of smooth potential functions that satisfy a symmetric Jacobian property [19], and Hurault et al. proposed the gradient step (GS) denoiser [32, 33], which can be interpreted as a proximal operator of an implicit regularizer [33]. Although the convergence of PnP-ADMM is commonly analyzed from the perspective of Douglas-Rachford splitting (DRS) based on their equivalence [49, 33], theoretically achieving fast fixed-point residual decay in PnP-DRS remains unknown.
Based on HPPP and denoiser priors,  [^1] Since the RED gradient exactly equals the residual, we adopt the GraRED notation—derived from Moreau decomposition—to represent the proximal operator of the implicit regularizer associated with the residual.we propose the Gradient Regularization by Denoising via HPPP called GraRED-HP3 for IR problems. The main contributions are as follows:
  1. 1.
    Theoretical Contributions. The sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP with the positive semi-definite preconditioner $\mathcal{M}$ converges strongly to a unique solution $\mathbf{u}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\|\mathbf{u}-\mathbf{a}\|_{\mathcal{M}}^{2}$, as stated in Theorem 3.1 and Proposition 3.3, while the original PPP method can only converge weakly to some uncertain solution. Moreover, compared with the asymptotic regularity result of PPP [9, Lemma 2.8] (i.e., $\lim_{k\to\infty}\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}=0$), we establish a convergence rate of $\mathcal{O}(1/k)$ for both $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|$ and $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}$ (see Propositions 3.10 and 3.12).
  2. 2.
    Algorithmic Development. We integrate HPPP with denoiser priors to propose the GraRED-HP3 algorithm for IR problems and shed new theoretical insights on denoiser priors from a PPP standpoint. Noting that the PnP-ADMM algorithm can be reformulated as an equivalent DRS form [49, 33], and given that the worst-case convergence rate for DRS is $\mathcal{O}(1/\sqrt{k})$ [28], GraRED-HP3 improved a sublinear convergence rate (as per Proposition 3.12) and thereby serves as an accelerated PnP method.
  3. 3.
    Experimental Validation. We numerically validate the advantages of HPPP, including its implicit regularity, acceleration, and efficiency with several toy examples. In addition, we demonstrate the state-of-the-art performance of GraRED-HP3 and its restarted variant with advanced denoisers in IR tasks.
The rest of this paper is organized as follows. In section 2, we review some preliminaries for convergence analysis. In section 3, we analyze the convergence of HPPP. Combining HPPP and denoiser priors, we propose GraRED-HP3 for IR problems. In section 4, we verify the advantages of HPPP with several toy examples. Furthermore, we validate the performance of GraRED-HP3 through IR experiments. Finally, conclusions are presented in section 5.
2 Preliminaries
[…]
2.1 Preconditioned proximal point
Bredies et al. [9] introduced the degenerate PPP framework with the positive semi-definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$. The proper preconditioner $\mathcal{M}$ can make $\mathcal{A}+\mathcal{M}$ have a lower triangular structure, which conveniently calculates the inverse $(\mathcal{A}+\mathcal{M})^{-1}$.
Definition 2.1.
An admissible preconditioner for the operator $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ is a bouned, linear, self-adjoint, and positive semi-definite operator $\mathcal{M}:\mathcal{H}\to\mathcal{H}$ such that
$\mathcal{T}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}$
is single-valued, full-domain, and Lipschitz continuous.
Therefore, the PPP iteration is written into
(6) $\mathbf{u}^{0}\in\mathcal{H},\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{M}+\mathcal{A})^{-1}\mathcal{M}\mathbf{u}^{k}.$
If $\mathcal{M}=I$, then $\mathcal{T}$ is firmly nonexpansive (FNE)operator [2, section 23] and (6) becomes the standard proximal point iteration. If $\mathcal{M}$ is positive semi-definite, the operator $\mathcal{T}$ is related with the degenerate $\mathcal{M}$-firmly nonexpansive ($\mathcal{M}$-FNE) operator [11, 15, 62, 9], which is associated with seminorm $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}$ and semi inner-product $\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{M}}=\langle\mathcal{M}\mathbf{u},\mathbf{v}\rangle$. The following notion extends monotone characteristics, i.e., $\mathcal{M}$-monotonicity.
Definition 2.2 ($\mathcal{M}$-monotonicity).
Let $\mathcal{M}:\mathcal{H}\to\mathcal{H}$ be a bouned linear positive semi-definite operator, then $\mathcal{B}:\mathcal{H}\to 2^{\mathcal{H}}$ is $\mathcal{M}$-monotone if we have
$\langle\mathbf{v}-\mathbf{v}^{\prime},\mathbf{u}-\mathbf{u}^{\prime}\rangle_{\mathcal{M}}\geq 0,\forall(\mathbf{u},\mathbf{v}),(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\in\mathcal{B}.$
The following lemma demonstrates that the firmly non-expansiveness notation can be generalized to the degenerate case [11, 15]. If $\mathcal{M}$ is an admissible preconditioner and $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone, then the operator $\mathcal{T}$ is $\mathcal{M}$-FNE.
Lemma 2.3 ([9]).
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone. Then $\mathcal{T}$ is $\mathcal{M}$-FNE, i.e.,
(7) $\left\|\mathcal{T}\mathbf{u}-\mathcal{T}\mathbf{v}\right\|_{\mathcal{M}}^{2}+\left\|(I-\mathcal{T})\mathbf{u}-(I-\mathcal{T})\mathbf{v}\right\|^{2}_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{v}\right\|^{2}_{\mathcal{M}}.$
By the KM iteration, the degenerate PPP method is written into
(8) $\mathbf{u}^{k+1}=(1-\lambda_{k})\mathbf{u}^{k}+\lambda_{k}\mathcal{T}\mathbf{u}^{k},$
where $\lambda_{k}$ is a sequence in $[0,2]$ such that $\sum_{k}\lambda_{k}(2-\lambda_{k})=+\infty$.
2.2 PnP and RED
2.2.1 PnP prior
Venkatakrishnan et al. proposed the first PnP method based on ADMM [56]. PnP methods replace the proximal mapping of the implicit regularizer defined by (9) with a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$. PnP-ADMM (Algorithm 1) is well known for its fast empirical convergence and efficiency in computational imaging [35, 60]. Sreehari et al. established theoretical conditions for PnP-ADMM, requiring that the Jacobian $\nabla D_{\sigma}$ be a doubly stochastic and symmetric matrix with all real eigenvalues in the range $(0,1]$ [51]. When the denoiser meets the necessary and sufficient conditions of Proposition 2.5, it is a proximal mapping of the implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$. Subsequently, a non-local means denoiser satisfying these conditions was constructed [51].
Algorithm 1 PnP-ADMM
1:Input: Given $\mathbf{u}^{0}=\mathbf{0},\mathbf{x}^{0},\mathbf{z}^{0}$ and total iterations $N>0$
2:for $k=0,1,2,\ldots,N-1$ do
3:     $\mathbf{x}^{k+1}=D_{\sigma}(\mathbf{z}^{k}-\mathbf{u}^{k})$
4:     $\mathbf{z}^{k+1}=\text{prox}_{\lambda f}(\mathbf{x}^{k+1}+\mathbf{u}^{k})$
5:     $\mathbf{u}^{k+1}=\mathbf{u}^{k}+(\mathbf{x}^{k+1}-\mathbf{z}^{k+1})$
6:end for
7:Output: $\mathbf{x}^{N}$
Definition 2.4 ([5]).
Given $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$, the proximal operator of $\phi(\mathbf{x})$ is defined by
(9) $\mathrm{prox}_{\phi}(\mathbf{x})=\arg\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{1}{2}\left\|\mathbf{u}-\mathbf{x}\right\|^{2}+\phi(\mathbf{x}).$
Proposition 2.5 ([40, 25]).
A function $h:\mathbb{R}^{n}\to\mathbb{R}^{n}$ defined everywhere is the proximal operator of a proper convex l.s.c. (lower semicontinuous) function $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ if, and only if the following conditions hold jointly:
  1. (a)
    there exists a convex l.s.c. function $\psi$ such that for each $\mathbf{x}\in\mathbb{R}^{n},h(\mathbf{x})=\nabla\psi(\mathbf{x})$;
  2. (b)
    $h$ is nonexpansive, i.e.,
    $\left\|h(\mathbf{x})-h(\mathbf{y})\right\|\leq\left\|\mathbf{x}-\mathbf{y}\right\|,\forall\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}.$
2.2.2 RED prior
Romano et al. introduced a famous RED [47], which is defined by
(10) $g_{\text{red}}(\mathbf{x})=\frac{1}{2}\langle\mathbf{x},\mathbf{x}-D_{\sigma}(\mathbf{x})\rangle,$
where $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a denoiser which is assumed to obey the following assumptions,
  1. (C1)
    Local Homogeneity: For all $\mathbf{x}\in\mathbb{R}^{n}$, $D_{\sigma}\left((1+\epsilon)x\right)=(1+\epsilon)D_{\sigma}(\mathbf{x})$ for sufficiently small $\epsilon>0$.
  2. (C2)
    Differentiability: The denoiser $D_{\sigma}(\cdot)$ is differentiable.
  3. (C3)
    Jacobian Symmetry [45]: $[\nabla D_{\sigma}(\mathbf{x})]^{\mathrm{T}}=\nabla D_{\sigma}(\mathbf{x}),\forall\mathbf{x}\in\mathbb{R}^{n}$.
  4. (C4)
    Strong Passivity: The spectral radius of the Jacobian satisfies $\eta\left(\nabla D_{\sigma}(\mathbf{x})\right)\leq 1$.
If $D_{\sigma}(\mathbf{x})$ satisfies C1C2, and C3, then $\nabla g_{\text{red}}(\mathbf{x})=\mathbf{x}-D_{\sigma}(\mathbf{x})=R(\mathbf{x})$. Moreover, if the denoiser satisfies C4, then RED is convex and $D_{\sigma}$ is nonexpansive.
2.2.3 Implicit Gradient RED
The gradient of RED is exactly the denoising residual. We show that under ideal conditions and assuming the residual is nonexpansive, both the denoiser and its residual function as proximal operators of implicit regularizers. Even if all RED conditions are hard to satisfy, learning FNE operators can still achieve this.
The nonexpansive assumption has theoretically played a crucial role in PnP and RED. Ryu et al. proposed the following assumption to guarantee the convergence of PnP methods [49],
(A) $\left\|(I-D_{\sigma})(\mathbf{u})-(I-D_{\sigma})(\mathbf{v})\right\|\leq\varepsilon\left\|\mathbf{u}-\mathbf{v}\right\|,$
for any $\mathbf{u},\mathbf{v}\in\mathbb{R}^{n}$ and $\varepsilon\leq 1$, the real spectral normalization is used to obtain nonexpansive residual. Under the conditons C1C2C3 and C4 of RED and assumption (A), we can further discuss the relationship between $\nabla g_{\text{red}}$ and the proximal operator, and prove that there exists an implicit regularizer $\phi$ such that $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$ and $R(\mathbf{x})=\nabla g_{\text{red}}(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, where $\phi^{*}$ is the conjugate of $\phi$.
Lemma 2.6.
Assume that a denoiser $D_{\sigma}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ satisfies conditons C1C2C3C4 and assumption (A), then there exists an implicit regularizer $\phi:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ such that
(11) $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x}).$
Proof 2.7.
By assumption condtions, the RED term $g_{\text{red}}$ is convex and $\nabla g_{\text{red}}=I-D_{\sigma}=R$. Let $\psi(\mathbf{x})=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$, then $D_{\sigma}=I-\nabla g_{\text{red}}=\nabla\psi$, which is a GS denoiser. Since $R$ is nonexpansive and $D_{\sigma}$ is differentiable, then $\eta(\nabla^{2}\psi)=\eta(\nabla D_{\sigma})=\eta(I-\nabla R)\in[0,1]$, it follows that $\psi$ is convex. Applying Proposition 2.5 together with $h(\mathbf{x})=D_{\sigma}(\mathbf{x}),\psi=\frac{1}{2}\left\|\mathbf{x}\right\|^{2}-g_{\text{red}}$ and Moreau decomposition [5, Theorem 6.44], i.e., $\mathrm{prox}_{\phi}(\mathbf{x})+\mathrm{prox}_{\phi^{*}}(\mathbf{x})=I$, then there exist a function $\phi$ and its conjugate $\phi^{*}$ such that $D_{\sigma}(\mathbf{x})=\nabla\psi(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x})$, and $R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$, which completes the proof.
Satisfying all RED conditions is challenging; in practice, only the nonexpansiveness assumption is required. From the perspective of monotone operator theory, Lemma 2.8 [2, Corollary 23.9] provides a necessary and sufficient condition for a denoiser to be a proximal operator of an implicit regularizer.
Lemma 2.8.
The mapping $T:\mathcal{X}\to\mathcal{X}$ is the resolvent of a maximal monotone operator if and only if $T$ is FNE.
Moreover, by [2, Proposition 4.4], the operator $T$ is FNE if and only if there exists a nonexpansive operator $S:\mathcal{X}\to\mathcal{X}$ such that $T=\frac{S+I}{2}$. Therefore, the core issue in explaining denoiser prior as a proximal operator of a convex regularizer is training a nonexpansive operator. Recent developments, such as spectral normalization [49, 21], learning maximally monotone operators [42], learning FNE denoisers [54, 10], and learning pseudo-contractive denoisers [57], have made the nonexpansive assumption more realistic in practice. Constructing the FNE denoiser $D_{\sigma}$ is feasible, and similarly, the residual $R=I-D_{\sigma}$ can also be FNE in (7) with $\mathcal{M}=I$. They can both be the proximal operators of implicit regularizers.
3 Halpern-type preconditioned proximal point (HPPP)
Compared with KM iteration, the classic Halpern iteration offers the advantage of strong convergence in infinite-dimensional Hilbert spaces [29]. The Halpern iteration [26]
$\mathbf{u}^{k+1}=\lambda_{k+1}\mathbf{u}^{0}+(1-\lambda_{k+1})T\mathbf{u}^{k}$
is an implicitly regularized method to find a particular fixed point [22], and the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by the Halpern iteration with suitable $\{\lambda_{k}\}_{k\in\mathbb{N}}$ strongly converges to the projection $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})$ [26, 37, 58, 30, 61, 44, 29], where $P_{\mathrm{Fix}(T)}(\mathbf{u}^{0})=\arg\min_{\mathrm{Fix}(T)}\left\|\mathbf{u}-\mathbf{u}^{0}\right\|^{2}$. When PPP is used to solve the inclusion problem (3), the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP can only converge weakly to some uncertain fixed point of $\mathcal{T}$. Based on the Halpern iteration, we propose the Halpern-type preconditioned proximal point (HPPP) algorithms (5).
3.1 Convergence analysis
[…]
Theorem 3.1.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}$ be the sequence generated by HPPP (5). Assume that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies
  1. (i)
    $\lim_{k\to\infty}\mu_{k}=0$;
  2. (ii)
    $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty$;
  3. (iii)
    $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0\text{ or }\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$.
Then $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ converges strongly to $\mathbf{u}^{*}$.
Proof 3.2.
First, we show $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$. Based on the definition of the upper limit, we take a subsequence $\{\mathbf{u}^{k_{n}}\}_{n\in\mathbb{N}}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ such that
(12) $\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}=\lim_{n\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}.$
By Lemma A.2(i), the sequence $\{\mathbf{u}^{k_{n}}\}_{n\in\mathbb{N}}$ is bounded, there exists a weakly convergent subsequence $\{\mathbf{u}^{k_{n_{j}}}\}_{j\in\mathbb{N}}$, we may assume $\mathbf{u}^{k_{n_{j}}}\rightharpoonup\tilde{\mathbf{u}}$. According to the conditions of the theorem, we have $\tilde{\mathbf{u}}\in\mathrm{Fix}(\mathcal{T})$.
Applying Lemma A.2 together with conditions (i)-(iii), we obtain $\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\to\mathbf{0}$, then $\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\rangle_{\mathcal{M}}\to 0$. Hence, we have
$\lim_{n\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$ $=\lim_{j\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},(\mathcal{T}\mathbf{u}^{k_{n_{j}}}-\mathbf{u}^{k_{n_{j}}})+(\mathbf{u}^{k_{n_{j}}}-\mathbf{u}^{*})\right\rangle_{\mathcal{M}}$
$=\lim_{j\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathbf{u}^{k_{n_{j}}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$
$=\left\langle\mathbf{a}-\mathbf{u}^{*},\tilde{\mathbf{u}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}.$
By Lemma B.1, we have
(13) $\limsup_{k\to\infty}\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}=\left\langle\mathbf{a}-\mathbf{u}^{*},\tilde{\mathbf{u}}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\leq 0.$
Denote $\delta_{k+1}=\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+2(1-\mu_{k+1})\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$, notice that
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*})\right\|_{\mathcal{M}}^{2}$
$=\mu_{k+1}^{2}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+(1-\mu_{k+1})^{2}\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}$
$+2\mu_{k+1}(1-\mu_{k+1})\left\langle\mathbf{a}-\mathbf{u}^{*},\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}$
$\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}^{2}+\mu_{k+1}\delta_{k+1},$
applying Lemma A.1 together with conditions (i)-(ii) and $\limsup_{k\to\infty}\delta_{k}\leq 0$, then we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0.$
By the definition of $\mathbf{u}^{k+1}$ we immediately obtain
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\|$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*})\right\|$
$\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+(1-\mu_{k+1})\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|$
(14) $\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}},$
where the last inequality follows from (26) and $C$ is defined in (25). Since $\lim_{k\to\infty}\mu_{k}=0$ and $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}=0$, and thus combined with (14) yields that
$\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|=0,$
which ends the proof of Theorem 3.1.
It is easy to verify that $\mu_{k}=\frac{1}{k^{\alpha}}(0<\alpha\leq 1)$ satisfes conditions (i)-(iii) of Theorem 3.1. Compared with [3], we have provided an alternative method to obtain the limit point as $\mathcal{M}$-projection of the initial point. The work [3] requires that $\mathcal{A}$ be linear, whereas we consider a more general case for $\mathcal{A}$. By introducing the maximal monotonicity assumption on $\mathcal{A}$, we can prove that every weak cluster point $\mathbf{u}$ of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$. According to Theorem 3.1, we obtain the following Proposition 3.3.
Proposition 3.3.
If the conditions of Theorem 3.1 hold, and $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ is a maximal monotone operator, then $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ converges strongly to $\mathbf{u}^{*}$.
Proof 3.4.
Assume that $\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. By Lemma A.2(iii), we have $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$, it follows that
$\mathcal{T}\mathbf{u}^{k_{n}}=(\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}})+\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{0}+\mathbf{u}=\mathbf{u},$
i.e., $\mathcal{T}\mathbf{u}^{k_{n}}\rightharpoonup\mathbf{u}$. Using $\mathcal{T}\mathbf{u}^{k_{n}}-\mathbf{u}^{k_{n}}\to\mathbf{0}$ we have
$\mathcal{A}\mathcal{T}\mathbf{u}^{k_{n}}\owns\mathcal{M}(\mathbf{u}^{k_{n}}-\mathcal{T}\mathbf{u}^{k_{n}})\to\mathbf{0}.$
Since $\mathcal{A}$ is maximal, we have that $\mathcal{A}$ is closed in $\mathcal{H}_{weak}\times\mathcal{H}_{strong}$ [2, Proposition 20.38], hence $\mathbf{0}\in\mathcal{A}\mathbf{u}$, i.e., $\mathbf{u}\in\mathrm{zer}\mathcal{A}=\mathrm{Fix}(\mathcal{T})$. Thus, we prove that every weak cluster point of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ lies in $\mathrm{Fix}(\mathcal{T})$. This completes the proof of Proposition 3.3.
Remark 3.5.
Compared Theorem 3.1, Proposition 3.3 with [9, Theorem 2.9] and [9, Corollary 2.10], where the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP converges strongly to a particular fixed point of $\mathcal{T}$. All conditions are the same except for additional assumption about $\{\mu_{k}\}_{k\in\mathbb{N}}$. The Lipschitz regularity of $(\mathcal{M}+\mathcal{A})^{-1}$ is a mild assumption especially in applications to splitting algorithms, which is used to prove the uniqueness of $\mathcal{M}$-projection and guarantee the boundedness of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$.
Hundal constructed an example in which the alternating projection $(P_{U}P_{V})^{k}\mathbf{x}^{0}$ converges weakly but not strongly, where $U$ is a closed convex cone and $V$ is a closed hyperplane [31]. Building on Hundal’s counterexample, [13] showed that the weak convergence of the DRS algorithm cannot be improved to strong convergence. Similarly, using the same counterexample, we can demonstrate that PPP also fails to converge strongly.
Counterexample 1.
Suppose that $\mathcal{H}$ is infinite-dimensional and separable. Let $V$ and $U$ be Hundal’s hyperplane and Hundal’s cone, respectively, and set
$\mathcal{A}_{1}:\mathbf{x}\mapsto\begin{cases}V^{\perp}&\text{ if }\mathbf{x}\in V,\\ \emptyset&\text{ if }\mathbf{x}\in V,\end{cases}\text{ and }\mathcal{A}_{2}=(P_{V}\circ P_{U}\circ P_{V})^{-1}-I.$
Then the sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by PPP (8) with $\lambda_{k}=1$ converges weakly, but not strongly, to a zero of $\mathcal{A}_{1}+\mathcal{A}_{2}$.
Proof 3.6.
As proved in [13], the operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are maximally monotone, and $\mathbf{0}\in\mathrm{zer}(\mathcal{A}_{1}+\mathcal{A}_{2})$, and the following sequence $\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$
(15) $\mathbf{w}^{k+1}=\mathbf{w}^{k}+J_{\mathcal{A}_{2}}(2J_{\mathcal{A}_{1}}\mathbf{w}^{k}-\mathbf{w}^{k})-J_{\mathcal{A}_{1}}\mathbf{w}^{k}$
generated by the DRS algortihm can weakly converges to $\mathbf{0}$ and $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$.
Since the above (15) can be viewed as a special PPP with
$\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&\mathcal{A}_{2}\end{pmatrix},\mathcal{M}=\begin{pmatrix}I&-I\\ -I&I\end{pmatrix},$
and $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k}$, where $\mathcal{C}^{*}=(I,-I)$. By [9, Corollary 2.15], $\mathbf{u}^{k}\rightharpoonup\mathbf{0}$. Indeed, if by contraction $\mathbf{u}^{k}\to\mathbf{0}$, then
$\left\|\mathbf{w}^{k}-\mathbf{0}\right\|=\left\|\mathcal{C}^{*}\mathbf{u}^{k}-\mathbf{0}\right\|\leq\left\|\mathcal{C}\right\|\cdot\left\|\mathbf{u}^{k}-\mathbf{0}\right\|\to 0,$
which contradicts the fact $\mathbf{w}^{k}\not\rightarrow\mathbf{0}$. Thus $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ can only converge weakly to $\mathbf{0}$.
Remark 3.7.
From Counterexample 1, HPPP has a strong convergence advantage over PPP in infinite-dimensional spaces, which has important theoretical value.
Remark 3.8.
Regarding the choice of the anchor point, we can use the degraded observed image, the denoised image, or the restarted technique (see Algorithm 2). The restarted technique is an effective method in machine learning [41] and LP [17, 38].
Algorithm 2 Restarted HPPP
1:Initialization $\mathbf{u}^{0,0}=\mathbf{a}^{0}$, total iteration $N>0$, period $q>0$, and epoch $N_{e}=\lfloor\frac{N}{q}\rfloor$
2:for $n=1,2,\cdots,N_{e}$ do
3:     for $k=0,1,\cdots,q-1$ do
4:         $\mathbf{u}^{n,k+1}=\mathrm{HPPP}(\mathbf{u}^{n,k},\mathbf{a}^{n})$
5:     end for
6:     $\mathbf{a}^{n+1}=\mathbf{u}^{n,q}$
7:end for
8:$\mathbf{u}^{N_{e},q}$
3.2 Convergence rate
Based on the following technical Lemma, Sabach and Shtern [50] first showed that, for a general form of Halpern iteration, the fixed-point residual converges at a rate of $\mathcal{O}(1/k)$.
Lemma 3.9 ([50]).
Let $M_{1}>0$. Assume that $\{a_{k}\}_{k\in\mathbb{N}}$ is a sequence of nonnegative real numbers such that $a_{1}<M_{1}$ and
$a_{k+1}\leq(1-\gamma b_{k+1})a_{k}+(b_{k}-b_{k+1})c_{k},k\geq 1,$
where $\gamma\in(0,1]$, the sequence $\{b_{k}\}_{k\in\mathbb{N}}$ is defined by $b_{k}=\min\{\frac{2}{\gamma k},1\}$, and $\{c_{k}\}_{k\in\mathbb{N}}$ is a sequence of real numbers such that $c_{k}\leq M_{1}<\infty$. Then the sequence $\{a_{k}\}_{k\in\mathbb{N}}$ satisfies
$a_{k}\leq\frac{M_{1}J}{\gamma k},k\geq 1,$
where $J=\lfloor\frac{2}{\gamma}\rfloor$.
By applying the Lemma 3.9, we will establish a sublinear convergence rate for $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|$ and fixed point residual $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|$ or $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$.
Proposition 3.10.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{u^{k}\}$ be the sequence generated by (5). If $\mu_{k}=\min\{\frac{2}{k},1\}$, then
  1. (i)
    $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$, i.e., $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}=\mathcal{O}(\frac{1}{k})$.
  2. (ii)
    $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$, $\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|=\mathcal{O}(\frac{1}{k})$.
Proof 3.11.
(i)According to (30), we have
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M\left|\mu_{k+1}-\mu_{k}\right|,$
where $M=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1},\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$, and $C_{1}$ is defined by (27). Applying Lemma 3.9 together with $c_{k}=M_{1}=M,J=2,\gamma=1,a_{k}=\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}$ and $\mu_{k}=b_{k}$ to the above inequality, we have $\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}\leq\frac{2M}{k}$.
(ii)By (33), when $k\geq 2$ we have
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ $\leq C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M^{\prime}\left|\mu_{k+1}-\mu_{k}\right|$
$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}=\mathcal{O}(\frac{1}{k}),$
where the second inequality follows from (i) and the fact that $\mu_{k}-\mu_{k+1}=\frac{2}{k(k+1)}(k\geq 2)$, and $M^{\prime}=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+CC_{1}$, and $C,C_{1}$ are defined in (25) and (27), respectively.
As for the second result, by the triangle inequality and $\mathbf{u}^{k+1}-\mathcal{T}\mathbf{u}^{k}=\mu_{k+1}(\mathbf{a}-\mathcal{T}\mathbf{u}^{k})$, for $k\geq 2$,
$\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|$ $\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\left\|\mathbf{u}^{k+1}-\mathcal{T}\mathbf{u}^{k}\right\|$
$\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|$
$\leq\frac{2MC}{k}+\frac{2M^{\prime}}{k(k+1)}+\frac{2M^{\prime}}{k+1}=\mathcal{O}(\frac{1}{k}).$
Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ according to [9, Proposition 2.3]. By [9, Theorem 2.13], the operator $\tilde{\mathcal{T}}=\mathcal{C}^{*}(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}$ is FNE, let $\mathbf{w}^{k}=\mathcal{C}^{*}\mathbf{u}^{k}$ and $\mathbf{a}=\mathbf{u}^{0}$, then HPPP is equivalent to the following reduced algorithm:
(16) $\mathbf{w}^{0}=\mathcal{C}^{*}\mathbf{u}^{0},\mathbf{w}^{k+1}=\mu_{k+1}\mathbf{w}^{0}+(1-\mu_{k+1})\tilde{\mathcal{T}}\mathbf{w}^{k}.$
By [36, Theorem 2.1] and [52, Proposition 2.9], we further give the following (tight) optimal convergence rate for HPPP.
Proposition 3.12.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be a maximal operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathbf{w}^{k}\}_{k\in\mathbb{N}}$ be the sequences generated by (5) and (16). If $\mu_{k}=\frac{1}{k+1}$ and $\mathbf{a}=\mathbf{u}^{0}$, then
  1. (i)
    $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|\leq\frac{2\left\|\mathbf{w}^{0}-\mathbf{w}^{*}\right\|}{k+1}$ for $k\geq 0$ and $\mathbf{w}^{*}\in\mathrm{Fix}(\tilde{\mathcal{T}})$.
  2. (ii)
    $\left\|\mathbf{u}^{k}-{\mathcal{T}}\mathbf{u}^{k}\right\|_{\mathcal{M}}\leq\frac{2\left\|\mathbf{u}^{0}-\mathbf{u}^{*}\right\|_{\mathcal{M}}}{k+1}$ for $k\geq 0$ and $\mathbf{u}^{*}\in\mathrm{Fix}({\mathcal{T}})$.
3.3 GraRED-HP3
The primal-dual algorithm to solve (4) is viewed as the fixed-point iteration $\mathbf{u}^{k+1}=\mathcal{T}\mathbf{u}^{k}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}\mathbf{u}^{k}$ with $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&\partial g^{*}\end{pmatrix},\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-\mathbf{K}^{*}\\ -\mathbf{K}&\frac{1}{s}I\end{pmatrix}$. Under the degenerate case $\tau s\left\|\mathbf{K}\right\|^{2}=1$, the HPPP iteration is given by
(17) $\left\{\begin{array}[]{ll}\mathbf{d}^{k}&=(I+\tau\lambda\partial f)^{-1}(\mathbf{x}^{k}-\tau\mathbf{K}^{*}\mathbf{y}^{k}),\\ \mathbf{x}^{k+1}&=\mu_{k+1}\mathbf{x}_{a}+(1-\mu_{k+1})\mathbf{d}^{k},\\ \mathbf{y}^{k+1}&=\mu_{k+1}\mathbf{y}_{a}+(1-\mu_{k+1})(I+s\partial g^{*})^{-1}\left(\mathbf{y}^{k}+s\mathbf{K}(2\mathbf{d}^{k}-\mathbf{x}^{k})\right),\end{array}\right.$
where $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a}),(\mathbf{x}_{0},\mathbf{y}_{0})\in\mathcal{X}\times\mathcal{Y}$ are the anchor and initial points. Under the ideal conditions of RED and with Assumption (A), according to Lemma 2.6, both the denoiser $D_{\sigma}$ and the residual $R$ can be expressed as the proximal operators of implicit regularizers, i.e., $D_{\sigma}(\mathbf{x})=\mathrm{prox}_{\phi}(\mathbf{x}),R(\mathbf{x})=\mathrm{prox}_{\phi^{*}}(\mathbf{x})$. Let $g(\mathbf{x})=\frac{\phi(s\mathbf{x})}{s}$, then $\phi^{*}=sg^{*}$ [5, Theorem 4.14], we can replace $(I+s\partial g^{*})^{-1}$ with the residual $R$ in (17). Set $\mathbf{K}=I$, the implicit Gradient RED via HPPP called GraRED-HP3 is proposed in Algorithm 3. Even if the conditions of RED are difficult to meet [45], recent developments [49, 54, 42, 21, 15, 10, 57] can learn FNE denoisers. According to Lemma 2.8, both the denoiser and the residual still act as the proximal operators of implicit regularizers.
3.3.1 Acceleration for PnP methods
PnP-ADMM (see Algorithm 1) is a well-known method to solve (1), which can be written into an equivalent DRS form [49, 33], i.e., $\mathbf{w}^{k+1}=\tilde{\mathcal{T}}\mathbf{w}^{k}=\mathbf{w}^{k}+D_{\sigma}(2\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})-\mathbf{w}^{k})-\mathrm{prox}_{\lambda f}(\mathbf{w}^{k})$. When $\mathbf{w}^{k}=\mathbf{x}^{k}-\mathbf{y}^{k},\tau=s=1,\mu_{k}=0$ in GraRED-HP3 results exactly in the DRS iteration. Thus, PnP-DRS is a special case of GraRED-HP3, which can be obtained from the perspective of PPP [11, 9]. The related GraRED-P3 algorithm is shown in Algorithm 4. Furthermore, Proposition 3.12 demonstrates that GraRED-HP3 is an accelerated PnP method achieving an $\mathcal{O}(1/k)$ convergence rate for the fixed-point residual.
Remark 3.13.
In [28], a worst-case convergence rate of $\mathcal{O}(\frac{1}{\sqrt{k}})$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$ was established for DRS, which also applies to PnP-ADMM. In contrast, GraRED-HP3 obtain an accelerated convergence rate of $\mathcal{O}(\frac{1}{k})$ for $\left\|\mathbf{w}^{k}-\tilde{\mathcal{T}}\mathbf{w}^{k}\right\|$.
Algorithm 3 GraRED-HP3
1:Initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, anchor point $\mathbf{a}=(\mathbf{x}_{a},\mathbf{y}_{a})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, and $R=I-D_{\sigma}$ is the residual.
2:for $k=0,1,2,\cdots,N-1$ do
3:     $\mathbf{d}^{k}=\mathrm{prox}_{\lambda f}(\mathbf{x}^{k}-\tau\mathbf{y}^{k})$
4:     $\mathbf{x}^{k+1}=\mu_{k+1}\mathbf{x}_{a}+(1-\mu_{k+1})\mathbf{d}^{k}$
5:     $\mathbf{y}^{k+1}=\mu_{k+1}\mathbf{y}_{a}+(1-\mu_{k+1})R\left(\mathbf{y}^{k}+s(2\mathbf{d}^{k}-\mathbf{x}^{k})\right)$
6:end for
7:$\mathbf{x}^{N}$.
Algorithm 4 GraRED-P3
1:Initialization $\mathbf{u}^{0}=(\mathbf{x}^{0},\mathbf{y}^{0})\in\mathbb{R}^{n}\times\mathbb{R}^{n}$, iteration number $N>0$, relaxing parameter $\lambda_{k}\in[0,2]$, and $R=I-D_{\sigma}$ is the residual.
2:for $k=0,1,2,\cdots,N-1$ do
3:     $\mathbf{d}^{k}=\mathrm{prox}_{\lambda f}(\mathbf{x}^{k}-\tau\mathbf{y}^{k})$
4:     $\mathbf{x}^{k+1}=\lambda_{k}\mathbf{d}^{k}+(1-\lambda_{k})\mathbf{x}^{k}$
5:     $\mathbf{y}^{k+1}=\lambda_{k}R(\mathbf{y}^{k}+s(2\mathbf{d}^{k}-\mathbf{x}^{k}))+(1-\lambda_{k})\mathbf{y}^{k}$
6:end for
7:$\mathbf{x}^{N}$.
Algorithm 5 Relaxed RED-PRO via SD
1:Input: $\mathbf{x}^{0}\in\mathbb{R}^{n}$, $\{t_{k}\}_{k\in\mathbb{N}}$, $\alpha\in(0,1)$, $\lambda$, $\mu$, $J$, $N,\delta>0$ and the denoiser $T$.
2:for $k=0,1,2,\ldots,N-1$ do
3:     $\mathbf{x}^{k,0}=\mathbf{x}^{k}$
4:     for $j=0,1,2,\ldots,J-1$ do
5:         $\mathbf{x}^{k,j+1}=T_{\alpha}\left(t_{j}\mathbf{x}^{k}+(1-t_{j})\mathbf{x}^{k,j}\right)$
6:     end for
7:     $\mathbf{v}^{k}=\frac{\delta}{\|\mathbf{x}^{k}-\mathbf{x}^{k,J}\|}\mathbf{x}^{k}+\left(1-\frac{\delta}{\|\mathbf{x}^{k}-\mathbf{x}^{k,J}\|}\right)\mathbf{x}^{k,J}$
8:     $\mathbf{x}^{k+1}=\mathbf{x}^{k}-\mu\left(\nabla f(\mathbf{x}^{k})+\lambda(\mathbf{x}^{k}-\mathbf{v}^{k})\right)$
9:end for
10:Output: $\mathbf{x}^{k+1}$
3.3.2 Compared with RED-PRO
Cohen et al. proposed the RED-PRO model [20],
(18) $\min_{\mathbf{x}\in\mathrm{Fix}(T)}\frac{1}{2\sigma^{2}}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|^{2}.$
RED-PRO sheds new insights on denoiser prior, which theoretically bridges between RED and PnP prior. However, practical denoisers often have narrow fixed-point sets, leading to suboptimal recovery solutions. To address this issue, they relax the hard constraint of $\mathrm{Fix}(T)$ and replace it with a dilated fixed-point set, defined for some $\delta>0$ as
$B_{\delta}(T)=\left\{\mathbf{x}\in\mathbb{R}^{n}:\left\|\mathbf{x}-P_{\mathrm{Fix}(T)}(\mathbf{x})\right\|\leq\delta\right\},$
where $P_{\mathrm{Fix}(T)}(\mathbf{x})$ is the fixed-point projection of $\mathbf{x}$ onto $\mathrm{Fix}(T)$. The relaxed RED-PRO (RRP) problem is formulated as
(19) $\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{1}{2\sigma^{2}}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|^{2}+\frac{\lambda}{2}\left\|\mathbf{x}-P_{B_{\delta}(T)}(\mathbf{x})\right\|^{2}.$
Algorithm 5 outlines RRP, where $T_{\alpha}=\alpha T+(1-\alpha)I$ with the demi-contractive denoiser $T$. When the steepest descent method solves (19), it requires an inner loop to calculate the fixed-point projection $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$, and the line 5 in RRP is equivalent to the Halpern iteration.
Comparing Algorithm 3 with the inner loop of Algorithm 5, both methods share the same algorithmic form of Halpern iteration and utilize identical anchor coefficients. In both cases, finding $P_{\mathrm{Fix}(T)}(\mathbf{x}^{k})$ or $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$ corresponds to a bilevel optimization problem that finds the point in the fixed-point set closest to the anchor point $\mathbf{x}^{k}$ or $\mathbf{a}$.
The key difference is that HPPP extends Halpern iteration to a degenerate form by employing a positive semi-definite metric $\mathcal{M}$, whereas the denoiser operator $T$ in RED-PRO is defined under the standard metric $\mathcal{M}=I$. From an optimization perspective, the RRP algorithm is based on gradient descent and uses the classic Halpern iteration in its inner loop. In contrast, $\mathcal{T}$ in GraRED-HP3 is an algorithmic operator and has no inner loop. From a PPP standpoint, we integrate HPPP with denoiser priors to propose Algorithm 3.
4 Experiments
In this section, we show the numerical experiments of the algorithms discussed in section 3. Firstly, we will verify the implicit regularity of HPPP by an easy 1D example, and validate the accelerated advantage of HPPP and Restarted HPPP by other toy examples. Then, we compare the CP, PPP (8), GraRED-HP3, and GraRED-P3 for image deblurring and inpainting under the same setting, and verify the efficiency of the proposed algorithms. The source code is available at https://github.com/zsc15/HPPP.
4.1 A toy example
We consider the following optimization problem in $\mathbb{R}$, i.e.,
(20) $\min_{x\in\mathbb{R}}f(x)+g(x),$
where $f(x)=\max\{-x,0\}$ and $g(x)=\max\{1-x,0\}$, we plot $f(x)+g(x)$ in Figure 1.
[…]
Figure 1: The image of $f(x)+g(x)$.
The corresponding saddle-point problem is $\min_{x\in\mathbb{R}}\max_{y\in\mathbb{R}}\{xy+f(x)-g^{*}(y)\}$, where $g^{*}(y)=\max_{x}[yx-\max\{1-x,0\}]=y+\delta_{[-1,0]}(y)$ and $\delta_{[-1,0]}$ is the indicator function of the interval $[-1,0]$. We denote the optimal set $X^{*}=[1,+\infty)=\arg\min_{x\mathbb{\in\mathbb{R}}}f(x)+g(x)$ and the primal-dual objective function $F(x,y)=xy+f(x)-g^{*}(y)$. Let us solve the saddle-point set $\{(x^{*},y^{*}):F(x^{*},y)\leq F(x^{*},y^{*})\leq F(x,y^{*}),\forall(x,y)\in\mathbb{R}^{2}\}$ of $F(x,y)$. Fixed $x^{*}\geq 1$, then
$\max_{-1\leq y\leq 0}\{x^{*}y-g^{*}(y)\}=\max_{-1\leq y\leq 0}\{x^{*}y-y\}=\left\{\begin{array}[]{ll}0,&x^{*}>1,y=0,\\ 0,&x^{*}=1,y\in[-1,0].\end{array}\right.$
If $x^{*}=1$, assume that $-1\leq y^{*}<0$, then $F(1,y^{*})=0$, while $F(x,y^{*})=y^{*}(x-1)+\max\{-x,0\}$ and there exists $x=2$ such that $F(2,y^{*})=y^{*}<F(1,y^{*})$, which leads to contradiction. Therefore, the saddle-point set is $\Omega=\{(x^{*},y^{*}):x^{*}\geq 1,y^{*}=0\}$.
When $\mathcal{M}=\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$, as shown in Figure 2, the minimization $\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\|\mathbf{u}-\mathbf{a}\|_{\mathcal{M}}^{2}=\arg\min_{(x,y)\in\Omega}(x-x_{a}-(y-y_{a}))^{2}$ can be interpreted as finding the $\mathcal{M}$-projection, namely, the projection of the anchor point $\mathbf{a}=(x_{a},y_{a})$ onto the line $x-y=x_{a}-y_{a}$. Therefore, the $\mathcal{M}$-projection of $\mathbf{a}$ onto $\Omega$ is uniquely determined,
(21) $P_{\Omega}^{\mathcal{M}}(\mathbf{a})=\begin{cases}(1,0),x_{a}-y_{a}-1\leq 0,\\ (x_{a}-y_{a},0),x_{a}-y_{a}-1>0.\end{cases}$
$x-y-1\leq 0$$x$$y$$O$$S(2,0)$$R(3,1)$
Figure 2: $\mathcal{M}$-projection onto the saddle point set.
We set the total iteration number $N=1000$. As shown in Figure 3(a), with the anchor point fixed at $\mathbf{a}=(12,9)$, the sequence $\mathbf{u}^{k}=(x^{k},y^{k})$ generated by HPPP converges to the same point $\mathbf{u}^{*}=(3,0)$, regardless of the initial point, which verifies Proposition 3.3 and (21). HPPP is also an implicitly regularized method. In contrast, Figure 3(b) shows that the sequence generated by PPP oscillates around the limit $(1.8,0)$. However, a quantitative mathematical characterization of the limit of the PPP sequence remains elusive.
Figure 3: Trajectory of $\mathbf{u}^{k}=(x^{k},y^{k})$ generated by HPPP or PPP.
4.2 A 2D toy example
We consider a $\frac{1}{\gamma}$-contractive operator $T_{\theta}:\mathbb{R}^{2}\to\mathbb{R}^{2}$ from [41], for any $\mathbf{x}=(x_{1},x_{2})^{\mathrm{T}}\in\mathbb{R}^{2}$,
$T_{\theta}\mathbf{x}=\frac{1}{\gamma}\begin{pmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{pmatrix}\mathbf{x}.$
We compare the Picard iteration $\mathbf{x}^{k+1}=T_{\theta}\mathbf{x}^{k}$, PPP ($\lambda_{k}=0.1$), HPPP ($\mu_{k}=\frac{1}{k+1}$) and Restarted HPPP ($q=50$). Set $\mathcal{A}=\begin{pmatrix}-T_{\theta}&I\\ -I&I\end{pmatrix}$ and $\mathcal{M}=\begin{pmatrix}\frac{1}{\tau}I&-I\\ -I&\frac{1}{s}I\end{pmatrix}$ with $\tau=0.8,s=1.25$, and initial points $\mathbf{x}^{0}=\mathbf{y}^{0}=\mathbf{x}_{a}=\mathbf{y}_{a}=(1,0)^{\mathrm{T}}$, where $\mathbf{a}=\mathbf{u}^{0}$. Figure 4 shows that HPPP and Restarted HPPP indeed provide accelerated residual decay.
Figure 4: Trajectories, distance to solution, two residuals of iterates for the 2D toy example. Here $\gamma=1/0.95,\theta=15^{\circ},N=100$.
4.3 An infinite-dimensional example
We consider an infinite dimensional problem on $\ell^{2}$ from [7], the space of square summable sequences. For any $\mathbf{x}=(x_{0},x_{1},x_{2},\cdots)\in\ell^{2}$, the norm is $\left\|\mathbf{x}\right\|=\left(\sum_{k\in\mathbb{N}}\left|x_{i}\right|^{2}\right)^{\frac{1}{2}}$. We consider the following operator inclusion problem:
(22) $\mathbf{0}\in\mathcal{A}_{1}\mathbf{x}+\mathcal{A}_{2}\mathbf{x},\mathcal{A}_{1}=-\mathcal{S}-\mathbf{b},\mathcal{A}_{2}=I,$
where $\mathcal{S}(\mathbf{x})=(0,x_{0},x_{1},\cdots)$ is the right-shift operator and $\mathbf{b}=(1,-1,0,\cdots)$. Since $\mathcal{A}_{1},\mathcal{A}_{2}$ are nonexpansive, then $\mathcal{A}_{1},\mathcal{A}_{2}$ are maximal monotone [2, Example 20.29]. Additionally, by solving the above inclusion problem (22) we obtain $\mathbf{x}^{*}=(1,0,\cdots)$. Using $\mathcal{A}=\begin{pmatrix}\mathcal{A}_{1}&I\\ -I&I\end{pmatrix}$ and positive semi-definite preconditioner $\mathcal{M}=\begin{pmatrix}2I&-I\\ -I&\frac{1}{2}I\end{pmatrix}$, we compare the PPP iteration with $\lambda_{k}=0.5,1,1.8$. We set $N=10^{3},\tau=0.5,s=2$ and initialize $\mathbf{x}^{0}=(0.9,1,0,\cdots)$ for HPPP and PPP, and the anchor point is the same as the initial point, i.e., $\mathbf{a}=(\mathbf{x}^{0},\mathbf{x}^{0})$. For this example, we present visual results about the residual and the distance to the solution. As illustrated in Figure 5, HPPP outperforms PPP in terms of both $\left\|\mathbf{x}^{k}-\mathbf{x}^{*}\right\|$ and $\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$.
(e) Distance to solution
(f) Residual comparison
Figure 5: Comparison with PPP to solve an infinite dimensional problem in $\ell^{2}$ space.
4.4 Image deblurring
Firstly, we compared with CP [16], PPP [9], and HPPP for TV-regularized IR problems. For TV-regularized IR problems,
(23) $\min_{\mathbf{x}\in\mathbb{R}^{n}}\frac{\lambda}{2}\left\|\mathbf{A}\mathbf{x}-\mathbf{y}\right\|_{2}^{2}+\beta\left\|\nabla\mathbf{x}\right\|_{1},$
where $\mathbf{y}$ is the degraded image, $\mathbf{A}$ is a linear operator, and $\lambda,\beta$ are balance coefficients.
We use a 2D Gaussian function with a standard deviation of 1.6 to convolve 10 test gray images, and finally obtain the degraded images with an additive white Gaussian noise with the noise level $\sigma=2.55$. Firstly, we compared three algorithms, CP, PPP, and the proposed HPPP. All the algorithms use the degraded images as initial points. We calculate the norm $K=\left\|\nabla\right\|=1.75$, and choose the total iteration $N=400$, balance coefficients $\lambda=2,\beta=5\times 10^{-4}$. Their parameters are given in Table 1. Both GraRED-P3 and GraRED-HP3 use DnCNN, other parameters are used below:
  • GraRED-P3 : $\tau=1,s=1,\lambda_{k}=0.2,\lambda=20$;
  • GraRED-HP3 : $\tau=1,s=1,\mu_{k}=1/(k+2),\lambda=20,\mathbf{x}_{a}=\mathbf{y},\mathbf{y}_{a}=\mathbf{0}$.
Table 2 shows PSNR (dB) of restoration results on CP, PPP, HPPP, GraRED-P3 and GraRED-HP3. The performance of these methods is evaluated using PSNR measure. The best recovery results are highlighted in bold. From Table 2, GraRED-P3 and GraRED-HP3 are better than classic algorithms with explicit TV regularization, which demonstrates implicit regularizer is more powerful to regularize inverse imaging problems. We visualize the numerical comparison between GraRED-P3, GraRED-HP3, CP, PPP, and HPPP in Figure 6. To further compare the robustness of the initial points between the proposed HPPP, CP, and PPP with TV regularization. As shown in Figure 7, we plot their respective evolutions of PSNR values for iterations for Butterfly and Parrots with 10 different random initial points, here we choose the degraded image $\mathbf{y}$ as the anchor point. Once the anchor point is chosen, the proposed HPPP algorithm is robust to random perturbations of the initial points, which achieves stable recovery.
Table 1: Parameters on CP, PPP, and HPP with TV regularization for image deblurring.
CP $\tau=s=1/K=0.57$
PPP $\tau=s=1/K=0.57,\lambda_{k}=1.95$ or $\lambda_{k}=1.2$
HPPP $\tau=s=1/K=0.57,\mu_{k}=\frac{1}{k+2},\mathbf{x}_{a}=\mathbf{A}^{\mathrm{T}}\mathbf{y},\mathbf{y}_{a}=0\cdot\nabla\mathbf{x}_{a}$
Table 2: Deblurring results of gray images compared with CP, PPP, and HPPP about the Gauss blurring kernel.
Cameraman House Pepper Starfish Butterfly Craft Parrots Barbara Boat
CP 26.04 30.86 25.99 27.53 27.85 25.51 26.88 24.46 28.99
PPP 26.05 30.85 25.99 27.53 27.85 25.51 26.88 24.46 28.98
HPPP 26.00 31.39 26.05 27.65 27.99 25.51 26.82 24.51 29.09
GraRED-P3 26.87 32.51 28.34 29.16 29.70 26.90 27.95 24.66 29.68
GraRED-HP3 26.85 32.51 28.80 29.15 29.67 27.03 27.89 24.66 29.99
Refer to captionRefer to caption
(a) Clean image
Refer to captionRefer to caption
(b) Degraded image
Refer to captionRefer to caption
(c) CP ($27.85$dB)
Refer to captionRefer to caption
(d) $\lambda_{k}=1.95$: PPP ($27.85$dB)
Refer to captionRefer to caption
(e) $\lambda_{k}=1.2$: PPP ($27.85$dB)
Refer to captionRefer to caption
(f) HPPP ($27.99$dB)
Refer to captionRefer to caption
(g) GraRED-P3 ($29.70$dB)
Refer to captionRefer to caption
(h) GraRED-HP3 ($29.67$dB)
Figure 6: Deblurring results of Butterfly degraded by the Gaussian blur kernel with noise level $\sigma=2.55$.
(a) Butterfly
(b) Parrots
Figure 7: Robustness comparison with random initialization. Test images are degraded by the Gaussian blurring kernel with the noise level $\sigma=2.55$.
Secondly, we compared with RED and RED-PRO. Following [47, 20], a $9\times 9$ uniform point spread function (PSF) or a 2D Gaussian function with a standard deviation of $1.6$ are used to convolve test images. We finally obtained the degraded images with the noise level $\sigma=\sqrt{2}$. The original RGB image is converted to the YCbCr image, all algorithms are applied to the luminance channel, and then the reconstruction image is returned to RGB space to obtain the final image. PSNR is measured on the luminance channel of the ground truth and the restored images. Table 3 shows the PSNR values of the restoration results on RED, RED-PRO, RRP, GraRED-P3, and GraRED-HP3. From the deblurring experiment Table 3 and Figure 9, GraRED-P3 and GraRED-HP3 achieve better performance than RED, RED-PRO, and RRP, which illustrates that KM or Halpern iteration used in PPP methods is effective. We further compare the differences among GraRED-P3, GraRED-HP3, and Restarted HPPP under six different settings with $\lambda=40,\tau=s=1,\lambda_{k}=0.2,\mu_{k}=\frac{1}{k+1},N=500,q=100$. In Table 5, we report the average PSNR values of 15 test images using six different blur kernels. When the anchor point and the initial point coincide, GraRED-HP3 achieves the best performance on some IR tasks. Moreover, by dynamically updating the anchor points, Restarted HPPP further boosts the performance of GraRED-HP3.
(a) Recovery peroformance
(b) Fixed point residual
Figure 8: Recovery peroformance and fixed-point residual comparison for the House image degraded by the Gaussian blurring kernel with the noise level $\sigma=5$. Here $D_{1}$ is used.
Additionally, we trained an FNE-DnCNN ($D_{1}$) using real spectral normalization [49], which is employed in RED and RED-PRO. We also compared with state-of-the-art methods using advanced denoisers such as DRUNet ($D_{2}$[65] and GS denoiser ($D_{3}$[32, 33, 59]. For GraRED-HP3, the anchor point is the same as the initial point. The default parameters are presented in Table 4. As shown in Table 6, Restarted HPPP with advanced denoisers achieves competitive or even superior performance compared to state-of-the-art methods. In Figure 8, we compare GraRED-P3 ($\lambda_{k}=0.1$) and GraRED-HP3 ($\mathbf{a}=\mathbf{u}^{0},\mu_{k}=\frac{1}{k+1}$). The results demonstrate that GraRED-HP3 achieves better recovery performance and exhibits a faster fixed-point convergence rate than GraRED-P3.
Table 3: Recovery results are obtained by different algorithms with TNRD [18]. The best two results are highlighted in red and blue colors, respectively.
Algorithms Uniform kernel Gaussian kernel($\sigma_{k}=1.6$)
Bike Butterfly Flower Hat Bike Butterfly Flower Hat
RED [47] 26.10 30.41 30.18 32.16 27.90 31.66 32.05 33.30
RED-PRO(HSD) [20] 24.95 27.24 29.38 31.55 27.36 30.55 31.81 33.07
RRP [20] 26.48 30.64 30.46 32.25 28.02 31.66 32.08 33.26
GraRED-P3 26.55 30.72 30.67 32.43 28.13 31.81 32.42 33.51
GraRED-HP3 26.80 30.88 30.74 32.42 28.06 31.80 32.27 33.50
Table 4: Parameters on GraRED-HP3 and Restarted HPPP with advanced denoisers.
[…]
Restarted HPPP$(D_{2})$ $\tau=1.2,s=1/1.2,N=30,q=10$
[…]
Table 5: Compared with GraRED-P3, GraRED-HP3, and Restarted HPPP with TNRD for different tasks.
kernels [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
noise level, scale $\sigma=5,\times 3$ $\sigma=2.55,\times 1$ $\sigma=2.55,\times 1$ $\sigma=2.55,\times 1$ $\sigma=2.55,\times 1$ $\sigma=2.55,\times 1$
GraRED-P3 27.49 29.25 30.92 28.19 33.53 33.73
GraRED-HP3 27.61 29.33 30.99 28.32 33.53 33.71
Restarted HPPP 27.66 29.36 31.00 28.34 33.53 33.73
Table 6: Compared with state-of-the-art methods. Test gray images are degraded by the Gaussian kernel with the noise level $\sigma=2.55$. The best two results are highlighted in red and blue colors, respectively.
Pepper Craft Cameraman Couple Man House Starfish Butterfly Boat
RED [47] 27.19 26.00 26.46 29.21 30.16 32.41 28.46 28.36 29.58
RED-PRO [20] 27.43 26.05 26.39 28.99 29.98 32.41 28.25 28.21 29.38
DPIR [65] 28.47 27.02 27.63 30.24 30.89 33.56 29.62 30.43 30.54
PnP-DRS [33] 29.79 26.92 27.46 29.70 29.59 33.25 29.94 30.66 29.46
PnP-PGD [32] 27.18 26.28 26.73 29.17 29.14 32.84 29.14 29.87 30.37
PnP-DYS [59] 27.22 26.27 26.62 28.97 29.04 32.87 29.06 29.88 28.68
[…]
Restarted HPPP$(D_{2})$ 28.34 26.75 27.02 29.71 30.44 32.69 28.91 29.28 30.01
[…]
Restarted HPPP$(D_{3})$ 29.81 27.13 27.49 30.04 30.80 33.15 30.09 30.53 30.37
Figure 9: Deblurring results of Butterfly degraded by the uniform kernel.
4.5 Image inpainting
In this section, we use the proposed algorithm to solve TV-regularized image inpainting problems and compare their numerical results with CP [16], PPP [9], and HPPP algorithms. The discrete image inpainting model is
$\min_{\mathbf{x}\in\mathbb{R}^{n}}\lambda\left\|\mathbf{M}\odot\mathbf{x}-\mathbf{y}\right\|_{F}^{2}+\beta\left\|\nabla\mathbf{x}\right\|_{1},$
where $\mathbf{y}$ is the degraded image, $\mathbf{M}$ is a mask, and $\lambda,\beta$ are balance coefficients.
We test 10 common images for evaluation. The first $\mathbf{M}$ is filled with a Bernoulli random mask whose each pixel is missing with probability $p=0.5$, i.e., $50\%$ of pixels are missed. The second $\mathbf{M}$ is a character mask where about $19\%$ of pixels are missed. All the algorithms start with the degraded image. For classic algorithms, we fix the balance parameter $\alpha=0.01$ and the total number $N=400$, and use the following other parameters:
  • HPPP: $\tau=s=1/\left\|K\right\|=0.57,\mathbf{x}_{a}=\mathbf{1}\in\mathbb{R}^{m\times n},\mathbf{y}_{a}=0\cdot\nabla\mathbf{x}_{a},\mu_{k}=\frac{1}{10(k+2)}$;
  • PPP: $\tau=s=1/\left\|K\right\|=0.57,\lambda_{k}=1.6$ or $\lambda_{k}=1.2$;
  • CP: $\tau=s=1/\left\|K\right\|=0.57$.
Both GraRED-P3 and GraRED-HP3 use DnCNN [49], other parameters are used below:
  • GraRED-P3 : $\tau=10,s=0.1,\lambda_{k}=0.2,\lambda=5$;
  • GraRED-HP3 : $\tau=10,s=0.1,\mu_{k}=0.05/(k+2),\lambda=5,\mathbf{x}_{a}=\mathbf{y},\mathbf{y}_{a}=\mathbf{0}$.
In Table 7 and Table 8, we compared the numerical performance of classic algorithms with TV regularization. In Figure 10 and 11, we compare visualization results of House degraded by the Bernoulli random mask and the character mask, the proposed algorithms achieve better visual performance than TV regularization.
Table 7: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=2.55$, Bernoulli random missing).
Cameraman House Peppers Starfish Butterfly Craft Parrots Barbara Boat
CP 23.54 28.94 24.53 23.89 23.59 23.42 23.20 23.10 26.29
PPP 23.58 29.05 24.53 24.32 23.76 23.43 23.33 23.13 26.36
HPPP 23.89 29.19 24.55 24.46 24.04 23.52 23.44 23.35 26.45
GraRED-P3 29.92 36.15 32.67 31.99 31.95 29.53 30.05 32.25 32.91
GraRED-HP3 29.90 36.14 32.68 31.97 31.92 29.52 30.04 32.23 32.90
Table 8: Numerical results of image inpainting compared with CP, PPP, and HPPP (noise level $\sigma=2.55$, character texture missing).
Cameraman House Peppers Starfish Butterfly Craft Parrots Barbara Boat
CP 25.87 31.58 29.40 26.16 25.60 26.33 25.20 24.21 25.23
PPP 26.33 31.78 30.13 26.86 26.08 26.47 25.64 27.32 28.05
HPPP 26.33 32.00 30.02 26.88 26.38 26.51 25.83 27.34 28.09
GraRED-P3 29.18 37.46 33.79 30.89 29.45 29.20 26.04 29.68 30.04
GraRED-HP3 29.13 37.47 33.80 30.93 29.42 29.20 26.02 29.60 30.04
Refer to captionRefer to caption
(a) Clean image
Refer to captionRefer to caption
(b) Degraded image
Refer to captionRefer to caption
(c) CP ($28.94$dB)
Refer to captionRefer to caption
(d) $\lambda_{k}=1.6$: PPP ($29.06$dB)
Refer to captionRefer to caption
(e) $\lambda_{k}=1.2$: PPP ($29.09$dB)
Refer to captionRefer to caption
(f) HPPP ($29.20$dB)
Refer to captionRefer to caption
(g) GraRED-P3 ($36.15$dB)
Refer to captionRefer to caption
(h) GraRED-HP3 ($36.14$dB)
Figure 10: Recovery results of House degraded by the random mask.
Refer to captionRefer to caption
(i) Clean image
Refer to captionRefer to caption
(j) Degraded image
Refer to captionRefer to caption
(k) CP ($31.58$dB)
Refer to captionRefer to caption
(l) $\lambda_{k}=1.6$: PPP ($31.79$dB)
Refer to captionRefer to caption
(m) $\lambda_{k}=1.2$: PPP ($31.74$dB)
Refer to captionRefer to caption
(n) HPPP ($32.00$dB)
Refer to captionRefer to caption
(o) GraRED-P3 ($37.46$dB)
Refer to captionRefer to caption
(p) GraRED-HP3 ($37.47$dB)
Figure 11: Recovery results of House degraded by the character mask.
5 Conclusions
This paper introduces a Halpern-type PPP (HPPP) algorithm, which has the advantage of strong convergence over PPP. In addition, GraRED-HP3 is proposed by integrating HPPP with denoiser priors for IR problems, which is an accelerated PnP-ADMM. Advantages of HPPP are demonstrated through several toy examples, and GraRED-HP3 achieves state-of-the-art performance on IR experiments. In the future, we plan to extend the definition of $\mathcal{M}$-monotonicity to $\mathcal{M}$-comonotonicity for nonconvex case [4].
Acknowledgments
The authors thank Prof. Defeng Sun for bringing the interesting work [52] to our attention and Mr. Feng Xue for his insightful discussions. We sincerely appreciate Associate Editor Michael Elad and the anonymous referees for their thorough review and valuable comments, which have substantially improved this paper.
Appendix A Boundedness
Firstly, we study the boundedness and asymptotic regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by (5). To further establish regularity of $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ (Theorem 3.1), we introduce a important lemma from [30, Lemma 2.5].
Lemma A.1.
Let $\{a_{k}\}_{k\in\mathbb{N}}$be a sequence of non-negative real numbers satisfying
(24) $a_{k+1}\leq(1-\mu_{k})a_{k}+\mu_{k}\beta_{k}+\gamma_{k},$
where $\{\mu_{k}\}_{k\in\mathbb{N}},\{\beta_{k}\}_{k\in\mathbb{N}},\{\gamma_{k}\}_{k\in\mathbb{N}}$ satisfies the following conditions:
  1. (i)
    $\{\mu_{k}\}$ converges to 0 in $[0,1]$, and $\sum_{k=0}^{\infty}\mu_{k}=+\infty$, or equivently $\prod_{k=0}^{\infty}(1-\mu_{k})=0$;
  2. (ii)
    $\lim\sup_{k\to\infty}\beta_{k}\leq 0$;
  3. (iii)
    $\gamma_{k}\geq 0,\sum_{k=0}^{\infty}\gamma_{k}<\infty$.
Then $\lim_{k\to\infty}a_{k}=0$.
Lemma A.2.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Let $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ be the sequence generated by (5). Then the following assertions are hold:
  1. (i)
    The sequences $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}},\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ are bounded;
  2. (ii)
    If the sequence $\{\mu_{k}\}_{k\in\mathbb{N}}$ satisfies $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$, then $\lim_{k\to\infty}\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}=0$;
  3. (iii)
    If the conditions of (ii) hold, then $\lim_{k\to\infty}\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=0,\lim_{k\to\infty}\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|=0$.
Proof A.3.
Let $\mathcal{M}=\mathcal{C}\mathcal{C}^{*}$ be a decomposition of $\mathcal{M}$ [9, Proposition 2.3]. Denote
(25) $C=L\left\|\mathcal{C}\right\|,$
for $\mathbf{u}^{\prime},\mathbf{u}^{\prime\prime}\in\mathcal{H}$, we have
$\left\|\mathcal{T}\mathbf{u}^{\prime}-\mathcal{T}\mathbf{u}^{\prime\prime}\right\|$ $=\left\|(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}\mathcal{C}^{*}\mathbf{u}^{\prime}-(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}\mathcal{C}^{*}\mathbf{u}^{\prime\prime}\right\|$
$=\left\|(\mathcal{M}+\mathcal{A})^{-1}\mathcal{C}(\mathcal{C}^{*}\mathbf{u}^{\prime}-\mathcal{C}^{*}\mathbf{u}^{\prime\prime})\right\|$
$\leq L\left\|\mathcal{C}\right\|\left\|\mathcal{C}^{*}(\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime})\right\|$
$=L\left\|\mathcal{C}\right\|\left\|\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime}\right\|_{\mathcal{M}}$
(26) $=C\left\|\mathbf{u}^{\prime}-\mathbf{u}^{\prime\prime}\right\|_{\mathcal{M}},$
where the third inequality follows from the $L$-Lipschitz property of $(\mathcal{M}+\mathcal{A})^{-1}$, and the fourth equality follows from the definition of $\left\|\mathbf{u}\right\|_{\mathcal{M}}$, i.e., $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}=\left\|\mathcal{C}^{*}\mathbf{u}\right\|$ for any $\mathbf{u}\in\mathcal{H}$.
First, we prove $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ is bounded. For any $\mathbf{u}^{*}\in\mathrm{Fix}(T)$, we have
$\left\lVert\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\rVert_{\mathcal{M}}$ $=\left\|\mu_{k+1}(\mathbf{a}-\mathbf{u}^{*})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*})\right\|_{\mathcal{M}}$
$\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}$
$\leq\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}},\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\}$
$\leq\dots\leq\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}},\left\|\mathbf{u}^{0}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\}.$
Denote
(27) $C_{1}=\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}},\left\|\mathbf{u}^{0}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\},$
then we have $\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\leq C_{1}$ for all $k\in\mathbb{N}$.
Furthermore,
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{*}\right\|$ $\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+(1-\mu_{k+1})\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|$
$\leq\mu_{k+1}\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+(1-\mu_{k+1})\cdot C\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}$
$\leq\max\{\left\|\mathbf{a}-\mathbf{u}^{*}\right\|,CC_{1}\}<+\infty.$
By (26), we have
(28) $\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{*}\right\|\leq C\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}\leq CC_{1}<+\infty.$
So the sequence $\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ is also bounded.
For (ii), we will prove $\lim_{k\to\infty}\left\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\right\|_{\mathcal{M}}=0$. Let $M=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1}$, by the triangle inequality and the $\mathcal{M}$-non-expansiveness of $\mathcal{T}$, we have
$\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+\left\|\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}$
$\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+\left\|\mathbf{u}^{k}-\mathbf{u}^{*}\right\|_{\mathcal{M}}$
(29) $\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|_{\mathcal{M}}+C_{1}=M.$
Note that $\mathbf{u}^{k+1}=\mu_{k+1}\mathbf{a}+(1-\mu_{k+1})\mathcal{T}\mathbf{u}^{k}$, we have
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|(\mu_{k+1}-\mu_{k})(\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1})+(1-\mu_{k+1})(\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1})\right\|_{\mathcal{M}}$
$\leq(1-\mu_{k+1})\left\|\mathcal{T}u^{k}-\mathcal{T}u^{k-1}\right\|_{\mathcal{M}}+\left|\mu_{k+1}-\mu_{k}\right|\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1}\right\|_{\mathcal{M}}$
(30) $\leq(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M\left|\mu_{k+1}-\mu_{k}\right|,$
where the last inequality follows from Lemma 2.3 and (29).
Applying Lemma A.1 together with $\lim_{k\to\infty}\frac{\mu_{k+1}-\mu_{k}}{\mu_{k}}=0$ or $\sum_{k\in\mathbb{N}}\left|\mu_{k+1}-\mu_{k}\right|<\infty$ to (30), we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|_{\mathcal{M}}=0$. Since $\mathbf{u}^{k+1}-\mathbf{u}^{k}=\mu_{k+1}(\mathbf{a}-\mathcal{T}\mathbf{u}^{k})$, we have
$\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}$ $=\left\|(\mathbf{u}^{k}-\mathbf{u}^{k+1})+(\mathbf{u}^{k+1}-\mathcal{T}{\mathbf{u}^{k}})\right\|_{\mathcal{M}}$
(31) $\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|_{\mathcal{M}}+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}.$
Using (29) and $\lim_{k\to\infty}\mu_{k}=0$ in (31), we can obtain $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|_{\mathcal{M}}=0$.
For (iii), let $M^{\prime}=\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+CC_{1}$, by the triangle inequality and (28), we have
(32) $\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+\left\|\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{a}-\mathbf{u}^{*}\right\|+CC_{1}=M^{\prime}.$
Similar to the proof of (30), we have
$\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|$ $\leq(1-\mu_{k+1})\left\|\mathcal{T}\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k-1}\right\|+\left|\mu_{k+1}-\mu_{k}\right|\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k-1}\right\|$
(33) $\leq C(1-\mu_{k+1})\left\|\mathbf{u}^{k}-\mathbf{u}^{k-1}\right\|_{\mathcal{M}}+M^{\prime}\left|\mu_{k+1}-\mu_{k}\right|,$
where the last inequality follows from (26). Applying (30) and $\lim_{k\to\infty}\mu_{k}=0$ to the above inequality (33), we have $\lim_{k\to\infty}\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=0$. Similar to the proof of (31),
$\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|\leq\left\|\mathbf{u}^{k}-\mathbf{u}^{k+1}\right\|+\mu_{k+1}\left\|\mathbf{a}-\mathcal{T}\mathbf{u}^{k}\right\|.$
By $\lim_{k\to\infty}\left\|\mathbf{u}^{k+1}-\mathbf{u}^{k}\right\|=0,\lim_{k\to\infty}\mu_{k}=0$ and boundedness of $\{\mathcal{T}\mathbf{u}^{k}\}_{k\in\mathbb{N}}$, we finally prove $\lim_{k\to\infty}\left\|\mathbf{u}^{k}-\mathcal{T}\mathbf{u}^{k}\right\|=0$.
Appendix B $\mathcal{M}$-projection
Lemma B.1.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Then there exists a unique solution ${\mathbf{u}}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ which solves:
$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T}).$
Proof B.2.
Since $l(\mathbf{u})=\frac{1}{2}\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}^{2}$ is a proper lower-semicontinuous differentiable convex function. Assume that $\mathbf{u}^{*}$ is the optimal solution of $\min_{\in\mathrm{Fix}(\mathcal{T})}l(\mathbf{u})$. Since $\mathrm{Fix}(\mathcal{T})=\mathrm{zer}\mathcal{A}$ is the convex set, thus $\mathbf{u}^{*}+t(\mathbf{u}-\mathbf{u}^{*})\in\mathrm{Fix}(\mathcal{T})$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$ and $t\in(0,1)$. Hence,
$\lim_{t\to 0}\frac{l(\mathbf{u}^{*}+t(\mathbf{u}-\mathbf{u}^{*}))-l(\mathbf{u}^{*})}{t}$ $=\left\langle l^{\prime}(\mathbf{u}^{*}),\mathbf{u}-\mathbf{u}^{*}\right\rangle$
$=\left\langle\mathcal{M}(\mathbf{u}^{*}-\mathbf{a}),\mathbf{u}-\mathbf{u}^{*}\right\rangle$
(34) $=\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0.$
If $\mathbf{u}^{**}$ is the another solution, then it also satisfies
(35) $\left\langle\mathbf{u}^{**}-{\mathbf{a}},\mathbf{u}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0.$
Replace $\mathbf{u}$ with $\mathbf{u}^{**},\mathbf{u}^{*}$ in (34) and (35), respectively, then
$\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}^{**}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,$
$\left\langle\mathbf{u}^{**}-\mathbf{a},\mathbf{u}^{*}-\mathbf{u}^{**}\right\rangle_{\mathcal{M}}\geq 0.$
Add above two inequalities, then we obtain $\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0$. By (26), we have
$\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|=\left\|\mathcal{T}\mathbf{u}^{*}-\mathcal{T}\mathbf{u}^{**}\right\|\leq C\left\|\mathbf{u}^{*}-\mathbf{u}^{**}\right\|_{\mathcal{M}}=0.$
It follows $\mathbf{u}^{*}=\mathbf{u}^{**}$, which completes the proof of Lemma B.1.
As mentioned in [3, 15], we can introduce the following notion of $\mathcal{M}$-projection.
Definition B.3 ($\mathcal{M}$-projection).
Assume $\mathbf{a}\in\mathcal{H}$, there exist an unique point $\mathbf{u}^{*}\in\mathrm{Fix}(\mathcal{T})$ such that $\left\|\mathbf{u}^{*}-\mathbf{a}\right\|_{\mathcal{M}}\leq\left\|\mathbf{u}-\mathbf{a}\right\|_{\mathcal{M}}$ for any $\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$, then $\mathbf{u}^{*}$ is called the $\mathcal{M}$-projection of $\mathbf{a}$ onto $\mathrm{Fix}(\mathcal{T})$, denoted by $P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$.
The following Lemma B.4 extends the variational inequality of the metric projection.
Lemma B.4.
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone and $(\mathcal{M}+\mathcal{A})^{-1}$ is $L$-Lipshitz. Then the following assertions are equivalent:
  1. (i)
    $\mathbf{u}^{*}=P_{\mathrm{Fix}(\mathcal{T})}^{\mathcal{M}}(\mathbf{a})$;
  2. (ii)
    $\left\langle\mathbf{u}^{*}-\mathbf{a},\mathbf{u}-\mathbf{u}^{*}\right\rangle_{\mathcal{M}}\geq 0,\forall\mathbf{u}\in\mathrm{Fix}(\mathcal{T})$.
Proof B.5.
See Lemma B.1 and Definition B.3.